What is pi(not pi/4) if we take the first 5 terms in this series?
pi/4=1-1/3+1/5-1/7+....

Answer by Fermat(127)

pi=4(1 - 1/3 + 1/5 - 1/7 + 1/9)
pi = 4(2/3 + 2/35 + 1/9)
pi = 4(263/315)
pi = 3 107/315
pi = 3.33968
============